14. Fix a non-zero vector i e R. Let L: R → R be the linear mapping defined by L(7) = – 2 proj„(7), for all a E R (a) Sho
Answer #1


L(x) = x-2x4-** [(y)=y-2x Porn [(y) = y eigen value is 1 yis orthogonal ton L(n) = n-2x***=- eigen valeus is (-1) (n-1) vecto